WAVE NUMBERS, STRENGTHS, WIDTHS AND SHIFTS WITH PRESSURE OF LINES OF ^{16}O_{2} IN SYS-TEMS a ^{1}Δ_{g} − X ^{3}Σ_{g}^{−} AND b ^{1}Σ_{g}^{+} − X ^{3}Σ_{g}^{−}

SWEE-LAN CHEAH, YUAN-PERN LEE, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; \underline{J. F. OGILVIE}, Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University drive, Burnaby, British Columbia V5A 1S6 Canada.